A ball is thrown straight upward at 36m/s from the ground level. How high will the...

Question 8 Unsaved

A ball is thrown straight upward at 36m/s from the ground level. How high will the ball go?
(Ignore the air friction)

Question 8 options:

38.9m

352.8m

66.1m

3.6m
Answer #1

u = 36 m/s

v = 0 m/s (since at the highest point, velocity is zero)

Use:

v^2 = u^2 – 2gS

=> S = (u^2 – v^2)/2g

= (36^2 – 0)/(2*9.8)

= 66.1 m